OpenStudy (anonymous):

Simplify the expression 3/(x+2) - 1/(3-x) - (2x-1)/(x^2-x-6)

OpenStudy (anonymous):

somehow I got the answer 2(x-4) over (x^2 +x +6)

OpenStudy (anonymous):

I got it by setting the three terms with a common denominator using the LCD I figured the LCD to be -(x-3)(x+2)

OpenStudy (anonymous):

what do you guys think?

OpenStudy (inkyvoyd):

That should be right.

OpenStudy (anonymous):

well I'm debating myself cuz my friend thinks otherwise. She says she keeps getting the answer: 2/x+2

OpenStudy (anonymous):

could you explain why it is 2/(x+2) ? I am confused...what did I do wrong?

OpenStudy (callisto):

3/(x+2) - 1/(3-x) - (2x-1)/(x^2-x-6) = 3/(x+2) + 1/(x-3) - (2x-1)/[(x-3)(x+2)] = [3(x-3)]/[(x-3)(x+2)] + 1(x+2)/[(x-3)(x+2)] -(2x-1)/[(x-3)(x+2)] = [(3x-9) + (x+2) - (2x-1)] / [(x-3)(x+2)] = (2x -6) /[(x-3)(x+2)] = 2(x-3) /[(x-3)(x+2)] = 2/(x+2)

OpenStudy (callisto):

you didn't show your work, how can we know what you did wrong ? :S

OpenStudy (anonymous):

^ you have a point Callisto :P